A Molecular Beam Optical Study of YCC

J. Xin, A. J. Marr and T. C. Steimle, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, 85287 -1604.

A weak signal near 13395 cm^{−1} in the low resolution laser induced fluorescence spectra of the laser ablated Y/ethylene supersonic expansion products was noted by Simard^{a}. He speculated that it may be due to yttrium dicarbide. Following up on his suggestion, we have detected by medium resolution LIF three red degraded bands at 12895 cm^{−1}, 13395 cm^{−1} and 13896 cm^{−1} which we attribute to YCC. The resolved LIF spectra for the 12895 cm^{−1} and 13395 cm^{−1} band systems produce strong Stokes progressions with the features separated by approximately 560 cm^{−1}. These are tentatively assigned to the Y-CC ground state stretching. Additional weaker features are also observed. The high resolution LIF spectrum of the 12895 cm^{−1} band system was recorded. Attempts to model the spectrum as a Ã^{4} A_{1}-X̃^{2}A_{1} transition suggested by a recent ab initio calculation^{b} are being made. Progress on the analysis will be presented.

aBenoit Simard, Steacie Institute-NRC, Ottawa Ontario, Canada, Private Communication.

bS. Roszak and K. Balasubramanian, Chem. Phys. Lett. 246, 20 (1995).